# Adjusted Connections I: Differential Cocycles for Principal Groupoid Bundles with Connection

### v1

@@ Abstract
We develop a new perspective on principal bundles with connection as morphisms from the tangent bundle of the underlying manifold to a classifying dg-Lie groupoid. This groupoid can be identified with a lift of the inner homomorphisms groupoid arising in Ševera’s differentiation procedure of Lie quasi-groupoids. Our new perspective readily extends to principal groupoid bundles, but requires an adjustment, an additional datum familiar from higher gauge theory. The resulting adjusted connections naturally provide a global formulation of the kinematical data of curved Yang–Mills–Higgs theories as described by Kotov–Strobl [KS15b] and Fischer [Fis21b].
@@ 1 Introduction and results
There are, however, more general situations, in which the matter fields do not take values in a vector space, but in a general manifold $M$, usually called the target space. If this manifold carries the action of a Lie group $\mathsf{G}\curvearrowright M$, then the scalar fields can be gauged. That is, the gauge and matter field content are combined into a connection on a principal bundle with the action groupoid $M\times\mathsf{G}\rightrightarrows M$ as structure groupoid. There is no other reason than technical simplicity for restricting to action groupoids, and the technology developed in this paper allows indeed for very general222and there are good reasons to believe that there is no more general formulation Lie groupoids as structure Lie groupoids. This leads directly to new forms of the famous Higgs mechanism that have so far been mathematically inaccessible and which may even have applications in phenomenology.
@@
Endowing such bundles with connections, however, is more subtle, even locally, i.e. over a coordinate patch of the base manifold over which the bundle trivializes. Consider, for example, the local description of connections as morphisms of differential graded (dg-)commutative algebras [Car50a, Car50b, KS15a] and in particular [SSS09]. For ordinary principal bundles, this description is obtained by restricting the global picture of a connection as a section of the Atiyah algebroid sequence [Ati57] to a patch $U$ of the base space. A connection is then a graded morphism $\mathcal{A}:\mathrm{T}[1]U\rightarrow\mathfrak{g}[1]$ from the grade-shifted tangent bundle to the grade-shifted structure Lie algebra [^3] See Section A for our conventions and more details on graded manifolds. Such morphisms are defined dually as morphisms of graded commutative algebras [^4] Recall that the Chevalley–Eilenberg algebra of a Lie algebra $\mathfrak{g}$ is the differential graded commutative algebra $\mathsf{CE}(\mathfrak{g})=\left.\bigodot\right.^{\bullet}\mathfrak{g}[1]^{*}$ with differential $\mathsf{d}a=-\tfrac{1}{2}[a,a]$ for $a\in\mathfrak{g}[1]^{*}$, see also Section 2.2. $\mathcal{A}^{\sharp}:\mathsf{CE}(\mathfrak{g})\rightarrow\Omega^{\bullet}(M)$. The curvature is then the failure of $\mathcal{A}^{\sharp}$ to respect the differential, and dg-commutative algebra morphisms correspond to flat connections.
This construction generalizes to higher Lie algebras as well as (higher) Lie algebroids $\mathfrak{g}$, but in many of these cases, the closure of the Lie algebra of gauge transformations restricts the curvatures strongly, and both these restrictions as well as the Bianchi identities arising in this construction contradict the expectations from physics, see e.g. [BJFJ+24] for a discussion.
@@
In the case where $\mathfrak{g}$ is concentrated in more than one degree, e.g. for $\mathfrak{g}$ a Lie algebroid or $\mathfrak{g}$ a higher Lie algebra, there is now an ambiguity in the identification of the generators that are mapped by $\mathcal{A}^{\sharp}$ to the additional higher connection forms. This is crucial, because different choices yield different definitions of curvatures, which induces different forms of gauge transformations. An adjustment is now such a choice with particularly nice properties. Importantly, the gauge transformations close without the above mentioned strong constraints on the curvatures, and in many cases, the curvatures gauge-transform in a nice, “covariant” manner, which is important for the construction of action principles for field theories involving them.
@@
So far, adjustments have been given mostly locally, and then for particular examples of higher Lie algebras, such as string-like higher Lie algebras [SSS09], their strict loop space models [SS20], twisted string and fivebrane structures [SSS12], and the higher Lie algebras appearing in the tensor hierarchies of gauged supergravity [BKS21]. Global descriptions have been given either abstractly, see e.g. [FSS14] for a weak string 2-group model, or explicitly, but only for particular examples, see [KS20, RSW22] for a strict string 2-group model; see also [BJFJ+24] for more details.
@@
In the above mention literature, the starting point was the construction of a gauge invariant action functional that provides a dynamical principle for a Lie-algebroid-valued connection. To this end, one has to define a covariantly transforming curvature, which is essentially achieved by adjusting the local connections as discussed above. This local adjustment is obtained from a connection on the Lie algebroid which satisfies several compatibility conditions, including a cohomological property of the connection and a certain curvature equation; flatness of the latter is essential for recovering ordinary gauge theory which is the reason why one speaks of a curved gauge theory.
@@ Results.
In this paper, we present the definition of $\mathsf{B}\mathscr{G}^{\rm conn}$ for $\mathscr{G}$ a Lie groupoid. We show that given a Cartan connection888see Section 2.3 for the definitions on the Lie groupoid $\mathscr{G}$, we can find an adjusted inner action groupoid $\mathscr{A}^{\mathsf{W}}(\mathscr{G})$, which takes the above mentioned role of $\mathsf{B}\mathscr{G}^{\rm conn}$.
We introduce three forms of adjustment: plain adjustments, covariant adjustments, and strict covariant adjustments, with the latter two requiring additional data. Covariant adjustments locally reproduces the notion of Lie algebroid-valued connections as developed in [KS15b] and [Fis21b].
We also present the details for a collection of interesting, concrete examples, focusing on fundamental Lie groupoids, Lie group bundles regarded as Lie groupoids with source and target both equal the projection, and action Lie groupoids. The last class contains all the examples known from the physics literature in the context of gauge-matter field theories. In particular, we present several examples which cannot be treated without introducing adjustments, such as Example 7.6, Example 7.12, and Example 7.25.
@@
It is also clearly interesting to explore the phenomenological implications and use cases for our construction in more detail. We will give a first concrete application in [FJKSc]. After this, it is clearly interesting to explore if some of the hopes voiced in [Str04] regarding phenomenological applications are justified.
@@ 2.1 Lie groupoids and their morphisms
A groupoid is a small category with invertible morphisms, and we write $\mathscr{G}=(\mathsf{G}\rightrightarrows M)$ with $M$ and $\mathsf{G}$ the set of objects and morphisms, respectively. A Lie groupoid is a groupoid $\mathscr{G}$ with $M$ and $\mathsf{G}$ smooth manifolds, where the source and target maps $\mathsf{s},\mathsf{t}\colon\mathsf{G}\rightrightarrows M$ are surjective submersions and the identity $\mathsf{e}\colon M\rightarrow\mathsf{G}$, and the composition $\circ\colon\mathsf{G}\mathbin{{}_{\mathsf{s}}\times_{\mathsf{t}}}\mathsf{G}\rightarrow\mathsf{G}$ are all smooth maps; as a consequence the inverse $-^{-1}\colon\mathsf{G}\rightarrow\mathsf{G}$ is then a smooth map, too. In the following, all manifolds are smooth.
@@
for $m,m_{1,2,3}\in M$.
@@
that respect morphism composition and identities in $\mathscr{G}$ and $\mathscr{H}$ and for which the left-action is principal, i.e. the map $\mathsf{G}\mathbin{{}_{\mathsf{s}}\times_{\tau}}B\rightarrow B\times_{N}B$, $(g,b)\mapsto(gb,b)$ is an isomorphism. Right-bibundles are defined analogously.
@@ 2.3 Connections on Lie groupoids and Lie algebroids
Concretely, we need the notion of Cartan connection on Lie groupoids, and we will follow closely [CSS12], but exchange $\ker(\mathrm{d}\mathsf{s})$ by $\ker(\mathrm{d}\mathsf{t})$; see also [AAC13] and in particular [Beh05].
@@
$L_{g}\colon\mathsf{G}^{\mathsf{s}(g)}\rightarrow\mathsf{G}^{\mathsf{t}(g)}~{}~{}~{}\mbox{and}~{}~{}~{}R_{g}\colon\mathsf{G}_{\mathsf{t}(g)}\rightarrow\mathsf{G}_{\mathsf{s}(g)}$ (2.11)
@@
This allows us to define the vector bundle isomorphisms [^12] Here and in the following, $\mathsf{p}_{M}$ always denotes the projection onto the base in some fiber bundle $E\rightarrow M$. Also, we are always very explicit in our notation of points in fiber bundles. This is to avoid confusion when the discussion gets more complicated later, where e.g. iterated tangent bundles appear.
$\mathscr{L}\colon\mathsf{G}\mathbin{{}_{\mathsf{s}}\times_{\mathsf{p}_{M}}}\mathfrak{g}$ $\rightarrow\ker(\mathrm{d}\mathsf{t})~{},~{}~{}~{}$ $\mathscr{R}\colon\mathsf{e}^{*}{\rm ker}(\mathrm{d}\mathsf{s})\mathbin{{}_{\mathsf{p}_{M}}\times_{\mathsf{t}}}\mathsf{G}$ $\rightarrow\ker(\mathrm{d}\mathsf{s})~{},$ (2.13)
$\bigl{(}g;\mathsf{s}(g),\nu\bigr{)}$ $\mapsto\mathrm{d}L_{g}\left(\mathsf{e}_{\mathsf{s}(g)},\nu\right)~{},~{}~{}~{}$ $\bigl{(}\mathsf{t}(g),\nu;g\bigr{)}$ $\mapsto\mathrm{d}R_{g}\left(\mathsf{e}_{\mathsf{t}(g)},\nu\right)~{},$
@@
$\mathrm{Ad}\colon\mathsf{s}^{*}\mathfrak{g}$ $\rightarrow\mathsf{t}^{*}\mathfrak{g}$ (2.18)
@@
where $\mathsf{inv}\colon\mathsf{G}\rightarrow\mathsf{G}$ is the groupoid inverse $\mathsf{inv}(g)=g^{-1}$, $\mathsf{inv}^{\star}$ is the isomorphism $\mathsf{inv}^{\star}\colon\mathsf{t}^{*}\mathfrak{g}\cong\mathsf{s}^{*}\mathfrak{g}$ induced by $\mathsf{inv}$.
@@
Remark 2.10 (Simplified notation).
Keeping track of the base point information in the pullback structures and of all the natural isomorphisms is most of the time irrelevant and implied by the context and the structure of the corresponding equation. In such a case one may simply write the adjoint action as
$\mathrm{Ad}(g,\nu)=-\left(\left.\vartheta^{\rm tot}\right|_{g^{-1}}\circ\mathrm{d}R_{g^{-1}}\right)\bigl{(}\mathrm{d}\,\mathsf{inv}(\nu)\bigr{)}=-\left(\mathrm{d}L_{g}\circ\mathsf{p}^{\rm vert}\circ\mathrm{d}R_{g^{-1}}\right)\bigl{(}\mathrm{d}\,\mathsf{inv}(\nu)\bigr{)}$ (2.22)
@@
$\mathrm{Ad}(g;\mathsf{s}(g),\nu)=\left(g;\mathsf{t}(g),\mathrm{d}L_{g}\left(\mathsf{e}_{\mathsf{s}(g)},\nu\right)\circ\sigma_{\mathcal{H}}\left(g^{-1};\mathsf{s}(g),\rho(\nu)\right)\right)$ (2.23)
@@
$\hskip 28.45274pt=\mathscr{L}\circ\mathsf{inv}^{\star}\circ\mathrm{Ad}(g,;\mathsf{s}(g),\nu)~{},$ (2.25)
@@
A connection on a groupoid $\mathscr{G}$ also induces a connection on the corresponding Lie algebroid $\mathfrak{g}$.
Proposition 2.12 ([CSS12, Thm. 3.11]).
Any Cartan connection $\mathcal{H}$ on a Lie groupoid $\mathscr{G}$ induces a connection on the Lie algebroid $\mathfrak{g}$ via
for all $\nu\in\Gamma(\mathfrak{g})$ and $V\in\mathfrak{X}(M)$, where $\nu_{l}$ is the left-invariant vector field on $\mathsf{G}$ induced by $\nu$, $\nu_{l}|_{g}=\mathscr{L}(g,\nu_{\mathsf{s}(g)})$. This Lie algebroid connection is Cartan, i.e. its basic curvature vanishes (see below).
Recall that a vector bundle connection on a Lie algebroid $\mathfrak{g}$ is a bilinear map
@@
such that the identities $\nabla_{fV}\nu=f\nabla_{V}\nu$ and $\nabla_{X}f\nu=f\nabla_{V}\nu+V(f)\nu$ holds for any smooth function $f$ on $M$.
@@
Let $\Theta\coloneqq\mathbbm{R}[-1]$. The tangent submersion $\sigma_{\Theta}(\mathcal{X})$ for $\mathcal{X}$ a manifold is the evident projection $\sigma_{\Theta}(\mathcal{X}):\mathcal{X}\times\Theta\rightarrow\mathcal{X}$. We write $\sigma_{\Theta}(*)$ for the trivial projection $\Theta\rightarrow*$, where * is the singleton set.
@@
The grade-shifted Lie algebroid $\mathsf{Lie}(\mathscr{G})[1]$ of a Lie groupoid $\mathscr{G}$ is the internal hom171717see Section A for details $\underline{\mathsf{hom}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ in the category of graded manifolds, which carries naturally the structure of a differential graded manifold. The differential in the corresponding Chevalley–Eilenberg algebra is induced by the canonical action of $\underline{\mathsf{hom}}(\Theta,\Theta)$ and the differential of $\mathscr{G}$.

### v3

@@ Abstract
We develop a new perspective on principal bundles with connection as morphisms from the tangent bundle of the underlying manifold to a classifying dg-Lie groupoid. This groupoid can be identified with a lift of the inner homomorphisms groupoid arising in Ševera’s differentiation procedure of Lie quasi-groupoids. Our new perspective readily extends to principal groupoid bundles, but requires an adjustment, an additional datum familiar from higher gauge theory. We show that for Lie groupoids, the additional adjustment data amounts to a Cartan connection. The resulting adjusted connections naturally provide a global formulation of the kinematical data of curved Yang–Mills–Higgs theories as described by Kotov–Strobl [KS15b] and Fischer [Fis21b].
@@ 1 Introduction and results
There are, however, more general situations, in which the matter fields do not take values in a vector space, but in a general manifold $M$, usually called the target space. If this manifold carries the action of a Lie group $\mathsf{G}\curvearrowright M$, then the matter fields can be gauged. That is, the gauge and matter field content are combined into a connection on a principal bundle with the action groupoid $M\times\mathsf{G}\rightrightarrows M$ as structure groupoid. There is no other reason than technical simplicity for restricting to action groupoids, and the technology developed in this paper allows indeed for very general222and there are good reasons to believe that there is no more general formulation Lie groupoids as structure Lie groupoids. This leads directly to new forms of the famous Higgs mechanism that have so far been mathematically inaccessible and which may even have applications in phenomenology.
@@
Endowing such bundles with connections, however, is more subtle, even locally, i.e. over a coordinate patch of the base manifold over which the bundle trivializes. Consider, for example, the local description of connections as morphisms of differential graded (dg-)commutative algebras [Car50a, Car50b, KS15a] and in particular [SSS09]. For ordinary principal bundles, this description is obtained by restricting the global picture of a connection as a section of the Atiyah algebroid sequence [Ati57] to a patch $U$ of the base space. A connection is then a graded morphism $\mathcal{A}\colon\mathrm{T}[1]U\rightarrow\mathfrak{g}[1]$ from the grade-shifted tangent bundle to the grade-shifted structure Lie algebra [^3] See Section A for our conventions and more details on graded manifolds. Such morphisms are defined dually as morphisms of graded commutative algebras [^4] Recall that the Chevalley–Eilenberg algebra of a Lie algebra $\mathfrak{g}$ is the differential graded commutative algebra $\mathsf{CE}(\mathfrak{g})=\left.\bigodot\right.^{\bullet}\mathfrak{g}[1]^{*}$ with differential $\mathsf{d}a=-\tfrac{1}{2}[a,a]$ for $a\in\mathfrak{g}[1]^{*}$, see also Section 2.2. $\mathcal{A}^{\sharp}\colon\mathsf{CE}(\mathfrak{g})\rightarrow\Omega^{\bullet}(M)$. The curvature is then the failure of $\mathcal{A}^{\sharp}$ to respect the differential, and dg-commutative algebra morphisms correspond to flat connections.
This construction generalizes to higher Lie algebras as well as (higher) Lie algebroids $\mathfrak{g}$, but in many of these cases, the closure of the Lie algebra of gauge transformations restricts the curvatures strongly, and both these restrictions as well as the Bianchi identities arising in this construction contradict the expectations from physics, see e.g. [BJFJ+25] for a discussion.
@@
In the case where $\mathfrak{g}$ is concentrated in more than one degree, e.g. for $\mathfrak{g}$ a Lie algebroid or $\mathfrak{g}$ a higher Lie algebra, there is now an ambiguity in the identification of the generators that are mapped by $\mathcal{A}^{\sharp}$ to the additional higher connection forms. This is crucial, because different choices yield different definitions of curvatures, which induce different forms of gauge transformations. An adjustment is now such a choice with particularly nice properties. Importantly, the gauge transformations close without the above mentioned strong constraints on the curvatures, and in many cases, the curvatures gauge-transform in a nice, “covariant” manner, which is important for the construction of action principles for field theories involving them.
@@
So far, adjustments have been given mostly locally, and then for particular examples of higher Lie algebras, such as string-like higher Lie algebras [SSS09], their strict loop space models [SS20], twisted string and fivebrane structures [SSS12], and the higher Lie algebras appearing in the tensor hierarchies of gauged supergravity [BKS21]. Global descriptions have been given either abstractly, see e.g. [FSS14] for a weak string 2-group model, or explicitly, but only for particular examples, see [KS20, RSW22] for a strict string 2-group model; see also [BJFJ+25] for more details.
@@
In the above mentioned literature, the starting point was the construction of a gauge invariant action functional that provides a dynamical principle for a Lie-algebroid-valued connection. To this end, one has to define a covariantly transforming curvature, which is essentially achieved by adjusting the local connections as discussed above. This local adjustment is obtained from a connection on the Lie algebroid which satisfies several compatibility conditions, including a cohomological property of the connection and a certain curvature equation; flatness of the latter is essential for recovering ordinary gauge theory which is the reason why one speaks of a curved gauge theory.
@@
We note that there are also constructions of Lie groupoid gauge theories without the use of adjustments [SS09]. Instead, a Moerdijk-Mrčun condition is imposed, i.e. the Ehresmann connection on the principal bundle as a horizontal distribution has to be in the kernel of what one calls the moment map $\Phi$ [^8] The action $p\mathbin{\vartriangleleft}g$ of a Lie groupoid element $g$ on a principal bundle element $p$ is defined if $\Phi(p)=\mathsf{t}(g)$, where $\mathsf{t}$ is the target arrow of the Lie groupoid. This condition, however, excludes a number of useful constructions. First, it excludes structural Lie group bundles as structural Lie groupoid as developed in [Fis22] because $\Phi$ is often the bundle projection of the principal bundle itself in such situations. In particular, we would lose many of the examples we present in Section 7. The Moerdijk–Mrčun condition would also exclude the previously-mentioned idea of covariantising Yang–Mills gauge theory. Covariantisation here includes two steps: first observing that a theory with a structural Lie group is equivalent to a theory with structural trivial Lie group bundle over the space-time (not just over a point) equipped with trivial connection, and second, extending the theory to non-trivial bundle structures, usually by extending the description to a richer family of connections, not just flat ones on the trivial group bundle. In particular, Yang–Mills gauge theory should be equivalent to a description with a structural trivial Lie group bundle as provided in [Fis22], but such a description is also excluded for the same reasons as mentioned above. Last but not least, our notion of Ehresmann connection generalises the one in [SS09], so that we effectively enlarge known gauge theory.
Furthermore, [CJ25] has a similar approach introducing an adjustment to define connections on a principal groupoid-bundle. However, in their context they look at an invariance w.r.t. representations up to homotopy, essentially avoiding the curvature condition observed in curved Yang-Mills-Higgs theories. In a future work one might investigate whether the work in [CJ25] is a kind of “curved Yang-Mills-Higgs theory up to homotopy”, similar to how representations up to homotopy are related to representations. (Observe that the same research team also works on higher gauge theory [CCK22]).
@@ Results.
In this paper, we present the definition of $\mathsf{B}\mathscr{G}^{\rm conn}$ for $\mathscr{G}$ a Lie groupoid. We show that the required adjustment datum, i.e. the additional algebraic information necessary for the general definition of a connection, is given by a Cartan connection999see Section 2.3 for the definitions on $\mathscr{G}$. Using this adjustment datum, we then construct an adjusted inner action groupoid $\mathscr{A}^{\mathsf{W}}(\mathscr{G})$, which takes the above mentioned role of $\mathsf{B}\mathscr{G}^{\rm conn}$.
We introduce three forms of adjustment: plain adjustments, covariant adjustments, and strict covariant adjustments, with the latter two requiring additional data. Covariant adjustments locally reproduces the notion of Lie algebroid-valued connections as developed in [KS15b] and [Fis21b]. As a new result at the local and infinitesimal level, we show that an adjustment here amounts to a Cartan connection on the relevant Lie algebroid.
We also present the a collection of detailed, interesting, and concrete examples, focusing on fundamental Lie groupoids, Lie group bundles regarded as Lie groupoids with source and target both equal the projection, and action Lie groupoids. The last class contains all the examples known from the physics literature in the context of gauge-matter field theories. In particular, we present several examples which cannot be treated without introducing adjustments, such as Example 7.6, Example 7.12, and Example 7.25.
@@
It is also clearly interesting to explore the phenomenological implications and use cases for our construction in more detail. We will give a first concrete application in [FJKSc]. After this, we plan to explore if some of the hopes voiced in [Str04] regarding phenomenological applications are justified.
@@ 2.1 Lie groupoids and their morphisms
A groupoid is a small category with invertible morphisms, and we write $\mathscr{G}=(\mathsf{G}\rightrightarrows M)$ with $M$ and $\mathsf{G}$ the set of objects and morphisms, respectively. A Lie groupoid is a groupoid $\mathscr{G}$ with $M$ and $\mathsf{G}$ smooth manifolds, where the source and target maps $\mathsf{s},\mathsf{t}\colon\mathsf{G}\rightrightarrows M$ are surjective submersions, and the identity $\mathsf{e}\colon M\rightarrow\mathsf{G}$ and the composition $\circ\colon\mathsf{G}\mathbin{{}_{\mathsf{s}}\times_{\mathsf{t}}}\mathsf{G}\rightarrow\mathsf{G}$ are all smooth maps; as a consequence, the inverse $-^{-1}\colon\mathsf{G}\rightarrow\mathsf{G}$ is then a smooth map, too. In the following, all manifolds are smooth.
@@
for $m,m_{1,2,3}\in M$. We have $\mathscr{P}\mathsf{air}(M)=\check{\mathscr{C}}(\sigma)$ for $\sigma=M\rightarrow*$.
@@
that respect morphism composition and identities in $\mathscr{G}$ and $\mathscr{H}$ and for which the left-action is principal, i.e. the map $\mathsf{G}\mathbin{{}_{\mathsf{s}}\times_{\tau}}B\rightarrow B\times_{N}B$, $(g,b)\mapsto(gb,b)$ is an isomorphism. Right-bibundles are defined analogously, cf., e.g., [Met03, Mrč96, Blo08].
@@ 2.3 Connections on Lie groupoids and Lie algebroids
Concretely, we need the notion of Cartan connection on Lie groupoids, and we will follow closely [CSS12], but exchange $\ker(\mathrm{d}\mathsf{s})$ by $\ker(\mathrm{d}\mathsf{t})$; the list of literature for Cartan connections and their properties is quite advanced, and historically, Cartan connections were not always known as Cartan connections, see [AAC13, Beh05, Bla06, Bla12, Bla16, CSS14, Tan06].
@@
$L_{g}\colon\mathsf{G}^{\mathsf{s}(g)}\rightarrow\mathsf{G}^{\mathsf{t}(g)}~{}~{}~{}\mbox{and}~{}~{}~{}R_{g}\colon\mathsf{G}_{\mathsf{t}(g)}\rightarrow\mathsf{G}_{\mathsf{s}(g)}~{},$ (2.11)
@@
This allows us to define the vector bundle isomorphisms [^13] Here and in the following, $\mathsf{p}_{M}$ always denotes the projection onto the base in some fiber bundle $E\rightarrow M$. Also, we are always very explicit in our notation of points in fiber bundles, and e.g. $(\mathsf{s}(g),\nu)$ is the point $\nu$ in the fiber over the base point $\mathsf{s}(g)$. This is to avoid confusion when the discussion gets more complicated later, where e.g. iterated tangent bundles appear.
$\mathscr{L}\colon\mathsf{s}^{*}\mathfrak{g}=\mathsf{G}\mathbin{{}_{\mathsf{s}}\times_{\mathsf{p}_{M}}}\mathfrak{g}$ $\rightarrow\ker(\mathrm{d}\mathsf{t})~{},$ (2.13)
$\bigl{(}g;\mathsf{s}(g),\nu\bigr{)}$ $\mapsto\mathrm{d}L_{g}\left(\mathsf{e}_{\mathsf{s}(g)},\nu\right)~{},$
and
$\mathscr{R}\colon\mathsf{t}^{*}\mathsf{e}^{*}{\rm ker}(\mathrm{d}\mathsf{s})=\mathsf{e}^{*}{\rm ker}(\mathrm{d}\mathsf{s})\mathbin{{}_{\mathsf{p}_{M}}\times_{\mathsf{t}}}\mathsf{G}$ $\rightarrow\ker(\mathrm{d}\mathsf{s})~{},$ (2.14)
$\bigl{(}\mathsf{t}(g),\nu;g\bigr{)}$ $\mapsto\mathrm{d}R_{g}\left(\mathsf{e}_{\mathsf{t}(g)},\nu\right)~{},$
@@
$\mathrm{Ad}_{\mathcal{H}}\colon\mathsf{s}^{*}\mathfrak{g}$ $\rightarrow\mathsf{t}^{*}\mathfrak{g}$ (2.19)
@@
where $\mathsf{inv}\colon\mathsf{G}\rightarrow\mathsf{G}$ is the groupoid inverse $\mathsf{inv}(g)=g^{-1}$, $\mathsf{inv}^{\star}$ is the isomorphism $\mathsf{inv}^{\star}\colon\mathsf{t}^{*}\mathfrak{g}\cong\mathsf{s}^{*}\mathfrak{g}$ induced by $\mathsf{inv}$. We may drop the subscript $\mathcal{H}$ if it is clear by context or if the adjoint representation is independent of the Cartan connection as in the case of Lie group bundles.
@@
$\mathrm{Ad}_{\mathcal{H}}(g;\mathsf{s}(g),\nu)=\left(g;\mathsf{t}(g),\mathrm{d}L_{g}\left(\mathsf{e}_{\mathsf{s}(g)},\nu\right)\circ\sigma_{\mathcal{H}}\left(g^{-1};\mathsf{s}(g),\rho(\nu)\right)\right)$ (2.23)
@@
$\hskip 28.45274pt=\mathscr{L}\circ\mathsf{inv}^{\star}\circ\mathrm{Ad}_{\mathcal{H}}(g,;\mathsf{s}(g),\nu)~{},$ (2.25)
@@
Remark 2.11 (Simplified notation).
Most of the time, the base point information in the pullback structures and of all the natural isomorphisms is implied by the context and the structure of the corresponding equation. In such a case one may simply write the adjoint action as
$\mathrm{Ad}_{\mathcal{H}}(g,\nu)=-\left(\left.\vartheta^{\rm tot}\right|_{g^{-1}}\circ\mathrm{d}R_{g^{-1}}\right)\bigl{(}\mathrm{d}\,\mathsf{inv}(\nu)\bigr{)}=-\left(\mathrm{d}L_{g}\circ\mathsf{p}^{\rm vert}\circ\mathrm{d}R_{g^{-1}}\right)\bigl{(}\mathrm{d}\,\mathsf{inv}(\nu)\bigr{)}$ (2.26)
A connection on a groupoid $\mathscr{G}$ also induces a vector bundle connection on the corresponding Lie algebroid $\mathfrak{g}$. Recall that a vector bundle connection on a Lie algebroid $\mathfrak{g}$ is a bilinear map
@@
such that the identities $\nabla_{fV}\nu=f\nabla_{V}\nu$ and $\nabla_{V}(f\nu)=f\nabla_{V}\nu+V(f)~{}\nu$ hold for any smooth function $f$ on $M$.
@@
Proposition 2.12 ([CSS12, Thm. 3.11]).
Any Cartan connection $\mathcal{H}$ on a Lie groupoid $\mathscr{G}$ induces a vector bundle connection on the Lie algebroid $\mathfrak{g}$ via
for all $\nu\in\Gamma(\mathfrak{g})$ and $V\in\mathfrak{X}(M)$, where $\nu_{l}$ is the left-invariant vector field on $\mathsf{G}$ induced by $\nu$, $\nu_{l}|_{g}=\mathscr{L}(g,\nu_{\mathsf{s}(g)})$. This connection is Cartan, i.e. its basic curvature vanishes.
Remark 2.13.
Analogous to [Fis22], one can express the invariance and the previous formulas also with the notion of the modified right-pushforward which may be more suitable for the dynamics in the related gauge theory.
@@
Let $\Theta\coloneqq\mathbbm{R}[-1]$. The tangent submersion $\sigma_{\Theta}(\mathcal{X})$ for $\mathcal{X}$ a dg-manifold is the evident projection $\sigma_{\Theta}(\mathcal{X}):\mathcal{X}\times\Theta\rightarrow\mathcal{X}$. We write $\sigma_{\Theta}(*)$ for the trivial projection $\Theta\rightarrow*$, where * is the singleton set.
@@
The grade-shifted Lie algebroid $\mathsf{Lie}(\mathscr{G})[1]$ of a Lie groupoid $\mathscr{G}$ is the internal hom181818see Section A for details $\underline{\mathsf{hom}}(\check{\mathscr{C}}(\sigma_{\Theta}(*)),\mathscr{G})$ in the category of graded manifolds, which carries naturally the structure of a differential graded manifold. The differential in the corresponding Chevalley–Eilenberg algebra is induced by the canonical action of $\underline{\mathsf{hom}}(\Theta,\Theta)$.